# Enhancing Multimodal Recommendations with Vision-Language Models and Information-Aware Fusion

### v1

@@ Abstract.
Recent advances in multimodal recommendation (MMR) have shown that incorporating rich content sources such as images and text can lead to significant gains representation quality. However, existing methods often rely on coarse visual features and uncontrolled fusion, leading to redundant or misaligned representations. As a result, visual encoders often fail to capture salient, item-relevant semantics, limiting their contribution in multimodal fusion. From an information-theoretic perspective, effective fusion should balance the unique, shared, and redundant information across modalities, preserving complementary cues while avoiding correlation bias. This paper presents VLIF, a vision-language and information-theoretic fusion framework that enhances multimodal recommendation through two key components. (i) A VLM-based visual enrichment module generates fine-grained, title-guided descriptions to transform product images into semantically aligned representations. (ii) An information-aware fusion module, inspired by Partial Information Decomposition (PID), disentangles redundant and synergistic signals across modalities for controlled integration. Experiments on three Amazon datasets demonstrate that VLIF consistently outperforms recent multimodal baselines and substantially strengthens the contribution of visual features [^1] Code is available at: https://github.com/dangkh/VLIF.
@@ 1. Introduction
Recently, multimodal recommendation (MMR) has emerged as a promising direction by incorporating diverse item-related content, including images, textual descriptions, and structured metadata. These additional sources offer rich semantic cues that complement behavioral signals, for instance, product images often emphasize features designed to attract buyers, providing valuable information.
@@
Figure 1. [Top] Our VLIF outperforms recent SOTA multimodal models. [Bottom] Using VLM to generate visual description of item.
@@
The second focuses on how to fuse heterogeneous modalities to construct expressive and semantically aligned representations (Xu et al., 2025a). Despite these advances, the visual modality has a limited contribution to overall performance (Zhou and Shen, 2023; Xu et al., 2025b), as removing image features typically results in only minor performance drops while textual signals remain dominant (Figure 1[Top]). In practice, product images are often optimized for marketing aesthetics rather than descriptive accuracy, which can mislead visual encoders and reduce the semantic fidelity of learned representations. This limitation highlights the need for incorporating textual context to guide visual understanding. Vision–Language Models (VLMs) offer a promising direction for enhancing multimodal recommendation by providing detailed, semantically aligned representations that capture fine-grained item characteristics and mitigate the limitations of raw visual features (Liu et al., 2024; Zhou et al., 2025). However, generic captioning or summarization often produces off-target content that fails to reflect the true semantics of the item. As illustrated in Figure 1[Bottom], when a product image contains multiple objects (e.g., a mother holding her daughter), a VLM prompted without item context may focus on irrelevant details, such as clothing rather than the actual baby carrier. Incorporating the item’s textual information (e.g., its title) allows the model to generate accurate, context-aware descriptions, demonstrating the value of text-guided visual enrichment in multimodal recommendation. Nevertheless, naive fusion strategies often treat modalities equally (Zhou et al., 2023; Xu et al., 2025b), introducing redundancy, conflicting cues, or over-reliance on text, which weakens representation quality. To overcome this, Partial Information Decomposition (PID) (Liang et al., 2023) provides a principled means to quantify modality interactions, allowing multimodal models to isolate complementary and consistent signals that enhance recommendation performance.
@@
To address these issues, we propose VLIF, a novel framework for multimodal recommendation. Instead of directly relying on raw image features, VLIF leverages vision–language models (VLMs) guided by item information (e.g., titles) to generate fine-grained, context-aware descriptions that capture essential item semantics. Furthermore, an information-aware fusion module inspired by PID is designed to quantify synergistic and redundant interactions between modalities, thereby enhancing multimodal representation learning. To summarize, the main contributions of this work are: First, we propose the first framework that leverage VLM to turn item images into enriched, meaningful features, reducing the limits of noisy or off-target representation. Second, we introduce a fusion module based on PID that strengthen the multimodal fusion. Third, extensive experiments on three Amazon product datasets demonstrate that VLIF consistently outperforms strong multimodal baselines.
Figure 2. Our VLIF pipeline
@@ 2. Methodology
We formulate our multimodal recommendation framework VLIF as follows. Let $\mathcal{U}$ and $\mathcal{I}$ denote the user and item sets. Each item $i\in\mathcal{I}$ is associated with a visual feature $\mathbf{v}_{i}$ and a textual feature $\mathbf{t}_{i}$, representing the visual and textual modalities, respectively.
As illustrated in Figure 2, our framework adopts a graph-based architecture that captures both heterogeneous and homogeneous relations to model multimodal user–item interactions. Each item embedding combines textual information with visual features, enhanced through a VLM-based Visual Enrichment module, which generates semantically aligned representations from item images. The resulting multimodal embeddings are further refined by an Information-Aware Fusion module for recommendation.
2.1. Multimodal Recommendation Framework
Previous studies (Zhang et al., 2021; Zhou et al., 2023) show that combining heterogeneous and homogeneous relations improves multimodal recommendation. Based on this insight, we build our framework upon a User–Item graph to model high-order collaborative relations across modalities.
@@ 2.1. Multimodal Recommendation Framework
Each modality-specific user and item embedding is then fed into the Information-Aware Fusion module (Section 2.3) to obtain unified representations $\mathbf{u}_{f}$ and $\mathbf{i}_{f}$, which are subsequently propagated through the homogeneous graphs to capture user–user and item–item relations.
Item–Item Graph. To model semantic correlations between items, we construct an item–item homogeneous graph via a $k$-nearest neighbor (KNN) strategy using item features from each modality $m$. The similarity between items $(i,i^{\prime})\in\mathcal{I}$ is measured by cosine similarity, and the final similarity score is defined as $S_{i,i^{\prime}}=\tfrac{1}{2}(S_{i,i^{\prime}}^{v}+S_{i,i^{\prime}}^{t})$, combining visual and textual similarities. Only the top-$k$ neighbors are retained, while the remaining edges are pruned. The item representation is updated as:
@@
where $\mathcal{N}_{i}$ denotes the neighbor set of item $i$ in the item–item graph and $\mathbf{h}_{i}^{(0)}$ is initialized with $\mathbf{i}_{f}$. For the User–User Graph, edges are constructed based on the number of co-interacted items, and only the top-$k$ neighbors are preserved. User representations are updated analogously to the aggregation process in Equation (3). The final user and item representations, $\mathbf{z}_{u}$ and $\mathbf{z}_{i}$, are obtained by integrating the homogeneous graph embeddings with the residual fusion signals. The predicted interaction score is then computed as:
@@ 2.2. VLM-based Visual Enrichment
To improve semantic reasoning and detail accuracy, we employ a Chain-of-Thought (CoT) prompting strategy that encourages step-by-step generation of enriched item descriptions. As illustrated in Figure 3, the VLM produces descriptive text, which is then encoded into a dense embedding serving as the enhanced visual representation:
@@ 2.3. Information-Aware Fusion
Cross-Modal Synergy Estimation (CSE). We first apply a cross-modal Transformer to capture interactive information between the two modalities:
@@
Transformer-based Redundancy Estimation (TBR). To extract the modality-shared (redundant) information $R$, we construct three input variants:
@@
This yields $\mathbf{v}^{\prime}$, the unique and redundancy-free visual representation orthogonal to $\mathbf{r}$. The final item representation concatenates the three complementary components—textual, synergistic, and unique visual features—as:
@@
where $||$ denotes concatenation. To model user-specific modality preferences, we introduce a learnable weight vector $\textbf{u}^{w}=[u^{w}_{0},u^{w}_{1},u^{w}_{2}]$ derived from user IDs, producing a fused user embedding:
@@ 2.4. Optimization
which maximizes the mutual information among $(h,h_{v},h_{t})$ and drives the model to capture a modality-invariant redundant representation $\mathbf{r}$. The final training objective is formulated as:
@@ 3.1. Experimental Settings
Dataset Metric BPR LightGCN VBPR MMGCN GRCN SLMRec LATTICE FREEDOM DRAGON MENTOR COHESION VLIF Improv.
@@
Dataset. Following prior studies (Xu et al., 2025b; Zhou and Shen, 2023), we conduct experiments on three categories of the Amazon Review dataset [^2] Publicly available at http://jmcauley.ucsd.edu/data/amazon/links.html : Baby, Sports and Outdoors, and Clothing, Shoes and Jewelry, denoted as Baby, Sports, and Clothing for brevity. We adopt the standard 5-core setting, ensuring that each user and item has at least five interactions.
Evaluation. We compare VLIF with the following SOTA methods: BPR, LightGCN (He et al., 2020), VBPR (He and McAuley, 2016), MMGCN (Wei et al., 2019), GRCN (Wei et al., 2020), SLMRec (Tao et al., 2023), LATTICE (Zhang et al., 2021), FREEDOM (Zhou and Shen, 2023), DRAGON (Zhou et al., 2023),MENTOR (Xu et al., 2025b), COHESION (Xu et al., 2025a). Performances are evaluated on 2 metrics: Recall@$10$ (R@1010) and NDCG@$10$ (N@1010). We follow the setting of  (Zhou and Shen, 2023) with a random data splitting of 8:1:1 for training, validation, and testing.
Implementation Details. We implement VLIF with Xavier initialization and a batch size of 1024. Hyperparameters are set with a learning rate of 0.001 and embedding dimension $d=64$. The model uses two layers for user–item, item–item, and user–user graphs, with top-$k=10$ and $\lambda=0.1$. Image descriptions are generated using the Qwen-VL 7B model. Experiments are conducted on an NVIDIA A5000 (24GB) GPU. VLM-generated descriptions and textual content (concatenated from each item’s brand, title, description, and category) are both encoded using Sentence-BERT [^3] https://huggingface.co/sentence-transformers.
@@ 3.2. Experimental Results
Overall Performance. As shown in Table 1, our VLIF consistently outperforms the two strongest baselines, MENTOR and COHESION, by a notable margin across all datasets and evaluation metrics. This demonstrates that integrating features derived from VLMs with our information-aware fusion module yields superior performance compared to using raw features or conventional fusion strategies, providing an effective solution for multimodal representation learning. When compared with (Zhou et al., 2023), which adopts a similar multimodal encoder, replacing the original visual features with VLM-based representations and our fusion strategy further improves performance.
Ablation. Furthermore, we investigate the contribution of each proposed module. Table 2 reports the performance when removing either the VLM-based feature generation or the information-aware fusion. The performance degradation observed highlights the necessity of both components. In particular, for image features, the original visual embeddings yield the lowest accuracy, followed by plain VLM features, while our task-specific VLM guided by item titles (VLIF) achieves the best performance. Likewise, alternative fusion strategies perform worse than our proposed information-aware fusion. Moreover, Figure 1 illustrates that incorporating VLM features balances the contribution of each modality, preventing over-reliance on textual signals. Finally, in Figure 4, we also evaluate VLIF with different VLM backbones, and the results confirm that our framework maintains stable and consistent improvements across models, demonstrating its general applicability. Furthermore, we visualize the original and VLM-based visual embeddings. The original features are scattered with many outliers, while the VLM-based embeddings show a more structured distribution, indicating richer and more semantically aligned representations.
@@ 4. Conclusion
In this work, we propose VLIF, a multimodal recommendation framework that leverages VLM to generate fine-grained item representations and an information-aware fusion module to integrate modalities effectively. Experiments on three Amazon product datasets demonstrate that VLIF consistently outperforms strong multimodal baselines and significantly strengthens the contribution of the visual modality. In future work, we plan to further enhance the framework by jointly training the VLM with recommendation signals, enabling task-adaptive visual–textual understanding for recommendation.

### v2

@@ Abstract.
Recent advances in multimodal recommendation (MMR) highlight the potential of integrating visual and textual content to enrich item representations. However, existing methods often rely on coarse visual features and naive fusion strategies, resulting in redundant or misaligned representations. From an information-theoretic perspective, effective fusion should balance unique, shared, and redundant modality information to preserve complementary cues. To this end, we propose VIRAL, a novel Vision–Language Information-aware Recommendation framework that enhances multimodal fusion through two components: (i) a VLM-based visual enrichment module that generates fine-grained, title-guided descriptions for semantically aligned image representations, and (ii) an information-aware fusion module inspired by Partial Information Decomposition (PID) to disentangle and integrate complementary signals. Experiments on three Amazon datasets show that VIRAL consistently outperforms strong multimodal baselines and substantially improves the contribution of visual features [^1] Code is available at: https://github.com/dangkh/VLIF.
@@ 1. Introduction
In recent years, multimodal recommendation (MMR) has emerged as a promising approach that integrates diverse item-related content—such as images, text, and metadata—to enrich behavioral signals. These additional modalities provide rich semantic cues that complement behavioral signals. For example, product images often highlighting features that attract buyers and enhance recommendation accuracy.
A key challenge in MMR is effectively fusing heterogeneous modalities to construct expressive, semantically aligned representations (COHESION). Despite notable progress, the visual modality contributes little to overall performance—removing image features typically causes only minor degradation, while textual signals remain dominant (FREEDOM; Mentor). This occurs because product images are often optimized for marketing aesthetics rather than descriptive accuracy, which can mislead visual encoders and reduce semantic fidelity (see Figure 1[Top]). These issues highlight the need to incorporate textual context to better guide visual understanding.
@@
Figure 1. [Top] Our VIRAL outperforms recent SOTA multimodal models. [Bottom] Using VLM to generate visual description of item.
@@
Vision–Language Models (VLMs) generate semantically rich representations that capture fine-grained item details and mitigate the limitations of raw visual features (Rec-gpt4v; largeVisionEmpirical). However, generic captioning often produces off-target content that fails to reflect true item semantics. As shown in  Figure 1[Bottom], when an image contains multiple objects (e.g., a mother holding her daughter), a VLM prompted without item context may focus on irrelevant details such as clothing rather than the actual product (e.g., a baby carrier). Incorporating textual cues (e.g., the item title) enables more accurate, context-aware descriptions, underscoring the value of text-guided visual enrichment in multimodal recommendation.
Nonetheless, naive fusion strategies (DRAGON; Mentor) often treat modalities equally, which can introduce redundancy, conflicting cues, or an over-reliance on textual information—ultimately degrading representation quality. To overcome these challenges, Partial Information Decomposition (PID) (quantifying) offers a principled framework for quantifying modality interactions. By isolating complementary and consistent information across modalities, PID-based approaches facilitate more effective fusion and lead to more robust, semantically aligned multimodal representations for recommendation tasks.
@@
To address these issues, we propose Vision–Language Information-aware Recommendation with Adaptive Learning (VIRAL), a novel framework for multimodal recommendation. Instead of directly relying on raw image features, VIRAL leverages VLMs guided by item information (e.g., titles) to generate fine-grained, context-aware descriptions that capture essential item semantics. Furthermore, an information-aware fusion module inspired by PID is designed to quantify synergistic and redundant interactions between modalities, thereby enhancing multimodal representation learning.
To summarize, our main contributions are threefold: (i) We propose VIRAL, a Vision–Language Information-aware Recommendation framework that leverages VLMs to generate semantically enriched visual features; (ii) We design an adaptive information-aware fusion module based on PID to balance complementary and redundant modality signals; (iii) Extensive experiments on three Amazon datasets show that VIRAL achieves superior performance, improved visual utilization, and enhanced interpretability over SOTA multimodal baselines.
Figure 2. Our VIRAL pipeline
@@ 2. Methodology
2.1. Problem and Solution Overview
Problem Formulation. We formulate the multimodal recommendation problem as predicting a user’s preference for items by jointly modeling behavioral and multimodal content signals. Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. Each user $u\in\mathcal{U}$ interacts with a subset of items $\mathcal{I}_{u}\subseteq\mathcal{I}$, forming an implicit feedback matrix $\mathbf{R}\in{0,1}^{|\mathcal{U}|\times|\mathcal{I}|}$. Each item $i\in\mathcal{I}$ is associated with a visual feature $\mathbf{v}_{i}$ and a textual feature $\mathbf{t}_{i}$, representing its visual and textual modalities. The goal is to learn a function $f(u,i)$ that estimates the likelihood of user $u$ interacting with item $i$ by effectively leveraging both behavioral and multimodal information.
Framework Overview. As illustrated in Figure 2, our proposed VIRAL framework operates in three main stages, each detailed in the following sections. First, adopts a graph-based architecture that captures both heterogeneous and homogeneous relations to model multimodal user–item interactions (Section 2.2). Next, a VLM-based Visual Enrichment module (Section 2.3) leverages VLM guided by item titles to generate fine-grained, semantically aligned visual representations, mitigating the noise of raw image features. Finally, an Information-Aware Fusion module (Section 2.4) inspired by PID disentangles redundant, synergistic, and unique modality information to produce unified multimodal embeddings for recommendation.
2.2. Graph-based Multimodal Interaction Modeling
To capture both behavioral and semantic dependencies in multimodal recommendation, VIRAL constructs three complementary graphs: a heterogeneous user–item interaction graph that models explicit behavioral relations, and two homogeneous graphs item–item and user–user to capture semantic correlations among items and users, respectively.
@@ 2.2. Graph-based Multimodal Interaction Modeling
Each modality-specific user and item embedding is then fed into the Information-Aware Fusion module (Section 2.4) to obtain unified representations $\mathbf{u}_{f}$ and $\mathbf{i}_{f}$, which are subsequently propagated through the homogeneous graphs to capture user, item respectively.
Item–Item Graph. To model semantic correlations between items, we construct an item–item homogeneous graph $\mathcal{G}_{i}$ via a $k$-nearest neighbor (KNN) strategy using item features from each modality $m$. The similarity between items $(i,i^{\prime})\in\mathcal{I}$ is measured by cosine similarity, and the final similarity score is defined as $S_{i,i^{\prime}}=\tfrac{1}{2}(S_{i,i^{\prime}}^{v}+S_{i,i^{\prime}}^{t})$, combining visual and textual similarities. Only the top-$k$ neighbors are retained, while the remaining edges are pruned. The item representation is updated as:
@@
where $\mathcal{N}_{i}$ denotes the neighbor set of item $i$ in $\mathcal{G}_{i}$ and $\mathbf{h}_{i}^{(0)}$ is initialized with $\mathbf{i}_{f}$. For the User–User Graph, edges are constructed based on the number of co-interacted items, and only the top-$k$ neighbors are preserved. User representations are updated analogously to the aggregation process in Equation (3). The final user and item representations, $\mathbf{z}_{u}$ and $\mathbf{z}_{i}$, are obtained by integrating the homogeneous graph embeddings with the residual fusion signals. The predicted interaction score is then computed as:
@@ 2.3. VLM-based Visual Enrichment
To improve semantic reasoning and accuracy, we employ a Chain-of-Thought (CoT) prompting strategy that encourages step-by-step generation of enriched item descriptions. As illustrated in Figure 3, the VLM produces descriptive text, which is then encoded into a dense embedding serving as the enhanced visual representation:
@@ 2.4. Information-Aware Fusion
Cross-Modal Synergy Estimation (CSE). We first utilize a cross-modal Transformer to model interactions between modalities:
@@
Transformer-based Redundancy Estimation (TBR). To extract shared (redundant) features $R$, we construct three input variants:
@@
This yields $\mathbf{v}^{\prime}$, the unique and redundancy-free visual representation orthogonal to $\mathbf{r}$. The final item representation concatenates three components—textual, synergistic, and unique visual features—as:
@@
where $||$ denotes concatenation. To model user-specific modality preferences, we employ a learnable weight vector $\textbf{u}^{w}=[u^{w}_{0},u^{w}_{1},u^{w}_{2}]$ derived from user IDs, producing a fused user embedding:
@@ 2.5. Optimization
which maximizes the mutual information among $(h,h_{v},h_{t})$ to learn a modality redundant feature $\mathbf{r}$. The final training objective is:
@@ 3.1. Experimental Settings
Dataset Metric BPR LightGCN VBPR MMGCN GRCN SLMRec LATTICE FREEDOM DRAGON MENTOR COHESION VIRAL Improv.
@@
Dataset. Following prior studies (Mentor; FREEDOM), we conduct experiments on three categories of the Amazon Review dataset [^2] Publicly available at http://jmcauley.ucsd.edu/data/amazon/links.html : Baby, Sports and Outdoors, and Clothing, Shoes and Jewelry, denoted as Baby, Sports, and Clothing for brevity. We adopt the standard 5-core setting, ensuring that each user and item has at least five interactions.
Evaluation. We compare VIRAL with the following SOTA methods: BPR, LightGCN (LightGCN), VBPR (VBPR), MMGCN (MMGCN), GRCN (RGCN), SLMRec (SLMREC), LATTICE (LATTICE), FREEDOM (FREEDOM), DRAGON (DRAGON),MENTOR (Mentor), COHESION (COHESION). Performances are evaluated on 2 metrics: Recall@$10$ (R@1010) and NDCG@$10$ (N@1010). We follow the setting of  (FREEDOM) with a random data splitting of 8:1:1 for training, validation, and testing.
Implementation Details. We implement VIRAL with Xavier initialization and a batch size of 1024, learning rate of 0.001, embedding dimension $d=64$ and two layers for user–item, item–item, and user–user graphs, with top-$k=10$ and $\lambda=0.1$. Image descriptions are generated using the Qwen-2.5-VL-7B model. Experiments are conducted on an NVIDIA A5000 (24GB) GPU. VLM-generated descriptions and textual content (concatenated from item’s brand, title, description, and category) are both encoded using Sentence-BERT [^3] https://huggingface.co/sentence-transformers.
@@ 3.2. Experimental Results
Overall Performance. As shown in Table 1, our VIRAL consistently outperforms the two strongest baselines, MENTOR and COHESION, by a notable margin across all datasets and evaluation metrics. This demonstrates that integrating features derived from VLMs with our information-aware fusion module yields superior performance compared to using raw features or conventional fusion strategies, providing an effective solution for multimodal representation learning. When compared with (DRAGON), which adopts a similar multimodal encoder, replacing the original visual features with VLM-based representations and our fusion strategy significantly improves performance.
Ablation. Furthermore, we investigate the contribution of each proposed module. Table 2 reports the performance when removing either the VLM-based feature generation or the information-aware fusion. The performance degradation observed highlights the necessity of both components. In particular, for image features, the original visual embeddings yield the lowest accuracy, followed by plain VLM features, while our task-specific VLM guided by item titles (VIRAL) achieves the best performance. Likewise, alternative fusion strategies perform worse than our proposed information-aware fusion. Moreover, Figure 1 illustrates that incorporating VLM features balances the contribution of each modality, preventing over-reliance on textual signals. Finally, in Figure 4, we also evaluate VIRAL with different VLM backbones, and the results confirm that our framework maintains stable and consistent improvements across models, demonstrating its general applicability. Furthermore, we visualize the original and VLM-based visual embeddings. The original features are scattered with many outliers, while the VLM-based embeddings show a more structured distribution, indicating richer and more semantically aligned representations.
@@ 4. Conclusion
In this work, we propose VIRAL, a multimodal recommendation framework that leverages VLM to generate fine-grained item representations and an information-aware fusion module to integrate modalities effectively. Experiments on three Amazon product datasets demonstrate that VIRAL consistently outperforms strong multimodal baselines and significantly strengthens the contribution of the visual modality. We plan to further enhance the framework by jointly training the VLM with recommendation signals, enabling task-adaptive visual–textual understanding for recommendation.